OpenStudy (stepht):

5 cos 2 θ = 1 for 0o < θ < 360o

OpenStudy (irishboy123):

\(5 cos 2 θ = 1\) so \( cos 2 θ = \frac{1}{5}\)  solve for \(2 \theta\) using you calculator, then for \( \theta\)